OpenStudy (anonymous):

which is bigger -2 thirds or -3 fourths

OpenStudy (anonymous):

-2/3 = -0.6666(6) -3/4 = -0.75 now draw the line  so the bigger one is -2/3 (it's the less negative one)

OpenStudy (anonymous):

Thank you, I'm new to this site.

OpenStudy (anonymous):

You're welcome, any doubt just ask, if you have any more doubts in this matter just draw the line on the paper, then you'll get used to it